OpenStudy (anonymous):

Find the area of the sector with a central angle of 60 degress and a radius of 5 inches. Round to the nearest tenth

OpenStudy (anonymous):

A.5.2 in square B.10.5in square C.13.1in square D.26.2insquare

OpenStudy (accessdenied):

A sector is just a fraction of the full circle. The area of a full circle is given by the formula: \(A = \pi r^2 \). If /we only have \(x^o\) of the circle, this is the fraction: \(x^o / 360^o \), and thus we set up the area of a sector: \( A = \dfrac{x^o}{360^o} \pi r^2 \) \(x^o = 60^o\), \(r = 5\)